OpenStudy (anonymous):

Solve for x. 5x - 11 = 7x - 3

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

5x-11=7x-3 -5x -5x -11=2x-3 +3 +3 -8=2x now just divide both sides by -8 and you will get x